From GlueXWiki
Jump to: navigation, search

August 15, 2013 FDC meeting


  1. Production
  2. Engineering [13], [14], [15] (Bill)
    • Survey results [16]
  3. Electronics (Chris, Nick)
  4. Tests at EEL126
  5. Other


Participants: Eugene, Bill, Nick, Simon, Vlad, Dave, Beni, Glenn, and Lubomir.


- On Wednesday (8/14) we moved the FDC from Blue Crab to EEl126 (see pictures attached). We installed tires on the cart at 30psi that worked as a cushion. Dave: had problems with the truck this morning that delayed the moving; still thinks will finish by Friday. If not, some stuff can be left at Blue Crab close to the doors to be moved on Monday. Eugene suggested to move the two granite tables left at Blue Crab (one went to Brian Cross) to the ESB temporary and then try to occupy some of "our space" in the TEDF or Test Lab. Tom is providing help for the transportation.

- End of last week we finished the cabling dry run: 1/4 of the signal cables were installed (see pictures attached). The additional rails for the mesh are causing some problems for the installation and service of the chambers being very close to the pre-amps and connectors. Especially, due to the four additional rails on the sides the density of the cables between the rails is very high so that it is difficult to work on the pre-amps behind them. It turned out that the axis of the mesh for this exercise was shifted with respect to the FDC axis and we had more space at the side where we did the cabling than we are supposed to. Other than that and the finding that we need more cuts on the skin for cable attachment, everything was successful and the dry run demonstrated that the current cabling scheme will work.


- We discussed the possibility to remove the four extra rails on the sides and to use the two main rails (for the detectors) to attach the mesh on the sides. We still need the top and bottom rails for the mesh. Bill: this is possible and will work on that. We need to make the upstream mesh bearing removable so that the CDC can be inserted inside the mesh. Bill thinks using the same rails for the detector and the mesh will not be a problem with respect to possible deformations.

- Bill showed several drawings (linked above) explaining the attachment of the pulling straps to the FDC bearings and the mesh: using plates and additional strap on the mesh. The same plates will be used to attach the grounding cables.

- Results of the FDC survey at Blue Crab are linked above. Four fiducials per package were installed but not spaced exactly by 90deg (+/-15). The package positions were adjusted (up-down, left-right) in order to have <0.5mm offset from the ideal position in X and Y. In Z we have deviations from the vertical position up to 0.75mm, but we could not estimate right away what fraction of the deviation is due to a tilt and what due to a deformation. For that, survey people were supposed to do flatness measurements, but they didn't. The final survey will be done in the Hall on the platform before the cabling. We all agreed it is not worth doing one more survey in the EEL.

- We discussed how the rails inside BCAL will be installed and how they will be adjusted to the rails on the three carts. Working inside BCAL to install/adjust the rails will not be easy, but there's no way around it. The carts as installed on the platform rails will be movable in all direction, so as soon as the distance between the main rails is OK (it has be adjusted/surveyed at Blue Crab) it will not be a problem. The mesh rails don't have to be adjusted very precisely.


- Nick fixed the 5 connectors on the signal cables. Thus the flat part on these cables gets shorter and Bill suggested connecting them to the pre-amps that don't require bending of the cable.

Tests at EEL126

- The FDC was connected to the gas Ar/CO2 40/60 at 100ccpm. The packages were connected in series i.e. there's some pressure change from the first to the last package. We have four independent bubblers (one per package) that are all bubbling. The pressure is 30-50Pa. The plan is to apply HV to all the packages and make sure all the HV channels are working.

- The spare package was connected to the Oxygen sensor showing 300ppm as of Thursday and 230ppm on the next day. The flow was increased from 50 to 220ccpm. There's more work on the signal/LV cabling and on the grounding to be done before we start testing the spare package with DAQ.

- Beni and Eugene discussed the possible fADC125 algorithms, a continuation from the tracking meeting.